# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 30 Mar 2026 v3

@@ 8 Concluding Remarks
remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity. This should not be confused with the classical support question for the distribution of $\varphi(n)/n$ over all integers, which remains open.
Appendix A Positive Density Near $1/2$
The following lemma gives an independent, Siegel–Walfisz-based proof that intervals near the endpoint $1/2$ carry positive limiting measure for $\varphi(p-1)/(p-1)$.
Lemma 11 (Positive density of primes with totient ratio near 1/2).
For every $\varepsilon>0$,
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}>0.$
Proof.
If $\varepsilon\geq 1$, then $\tfrac{1}{2}-\varepsilon\leq 0$, so the displayed set contains every prime and the claim is trivial. Hence assume throughout that $0<\varepsilon<1$.
Step 1: Sieving out small odd prime divisors. Let $R\geq 3$ be a fixed integer, to be chosen later in terms of $\varepsilon$, and define
$M_{R}:=2\!\!\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}r,\qquad\nu_{R}:=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}(r-2),\qquad\delta(R):=\frac{\nu_{R}}{\varphi(M_{R})}=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}\frac{r-2}{r-1}.$
Also set
[…]
The finitely many odd primes $p\leq R$ contribute only $O_{R}(1)$ to $|\mathcal{A}_{R}(x)|$. Hence, by Theorem 5,
$|\mathcal{A}_{R}(x)|=\frac{\nu_{R}}{\varphi(M_{R})}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\delta(R)\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr).$
Applying Theorem 5 also with modulus $1$, we have
$\pi(x)=\operatorname{Li}(x)+O\!\bigl(xe^{-c\sqrt{\log x}}\bigr),$
and therefore
$\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}\longrightarrow\delta(R)>0.$
In particular, $\mathcal{A}_{R}(x)$ has positive relative density $\delta(R)$ among the primes.
Step 2: Reduction to a mean-value estimate. For $p\in\mathcal{A}_{R}(x)$, the only prime $\leq R$ dividing p1p-1 is $2$, so
$\frac{\varphi(p-1)}{p-1}=\frac{1}{2}\prod_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr).$
Set
$S(p):=\sum_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\frac{1}{q}.$
By the elementary inequality
$\prod_{i}(1-a_{i})\geq 1-\sum_{i}a_{i}\qquad(0\leq a_{i}\leq 1),$
any $p\in\mathcal{A}_{R}(x)$ with $S(p)\leq\varepsilon$ satisfies
$\frac{\varphi(p-1)}{p-1}\geq\frac{1}{2}\bigl(1-S(p)\bigr)\geq\frac{1}{2}(1-\varepsilon)>\frac{1}{2}-\varepsilon.$
It therefore suffices to show that a positive proportion of primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$.
Step 3: Mean-value estimate for $S(p)$. Throughout Steps 3 and 4, the parameter $R$ (and hence $\varepsilon$) is fixed; all limits are as $x\to\infty$. Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
$S_{\leq Q}(p)=\sum_{\begin{subarray}{c}q\mid(p-1)\\ R<q\leq Q\end{subarray}}\frac{1}{q}.$
For the tail, each prime divisor $q>Q$ of p1p-1 contributes at most $1/Q$, and there are at most $\log p/\log Q$ such divisors. Hence, uniformly for $p\leq x$,
$S_{>Q}(p)\leq\frac{\log p}{Q\log Q}\leq\frac{\log x}{Q\log Q}=\frac{1}{2\log x\,\log\log x}=o(1).$
Now assume $x$ is large enough that $Q>R$. For each prime $q\in(R,Q]$, write
$B_{q}(x):=\#\{p\in\mathcal{A}_{R}(x):p\equiv 1\pmod{q}\}$.
Since $q>R$, we have $q\nmid M_{R}$. The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is exactly a single admissible class, and modulo $q$ the residue class is forced to be $1$. Because $\varphi(M_{R}q)=\varphi(M_{R})(q-1)$, we obtain from Theorem 5 that
$B_{q}(x)=\frac{\nu_{R}}{\varphi(M_{R}q)}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\frac{\delta(R)}{q-1}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr),$
uniformly for $R<q\leq Q$. Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.
Dividing by the asymptotic formula for $|\mathcal{A}_{R}(x)|$ yields
$\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}=\frac{1}{q-1}+O_{R}\!\bigl((\log x)e^{-c_{R}\sqrt{\log x}}\bigr)$
uniformly for $R<q\leq Q$. Therefore
$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S_{\leq Q}(p)$ $=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q}\,\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}$
$\leq\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+O_{R}\!\Bigl((\log x)e^{-c_{R}\sqrt{\log x}}\sum_{q\leq Q}\frac{1}{q}\Bigr)$
$=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+o(1),$
since $\sum_{q\leq Q}1/q\ll\log\log Q\ll\log\log\log x$.
Combining this with the uniform bound on $S_{>Q}(p)$, we get
$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S(p)\leq\eta(R)+o(1),\qquad\eta(R):=\sum_{\begin{subarray}{c}q>R\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}.$
Since $\sum_{q}1/(q(q-1))$ converges, we have $\eta(R)\to 0$ as $R\to\infty$.
Step 4: Markov’s inequality and conclusion. Choose $R$ large enough that $\eta(R)<\varepsilon/2$. By Markov’s inequality, the fraction of primes in $\mathcal{A}_{R}(x)$ with $S(p)>\varepsilon$ is at most
$\frac{\eta(R)+o(1)}{\varepsilon}<\frac{1}{2}$
for all sufficiently large $x$. Hence at least half the primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$, and therefore
$\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon.$
It follows that for all sufficiently large $x$,
$\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}\geq\frac{1}{2}\,|\mathcal{A}_{R}(x)|.$
Dividing by $\pi(x)$ and taking $\liminf$, we conclude that
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}$ $\geq\frac{1}{2}\lim_{x\to\infty}\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}$
$=\frac{\delta(R)}{2}>0.$
This proves the lemma. ∎

### 21 Apr 2026 v4

@@ 8 Concluding Remarks
remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity. The numerical evidence of Remark 2 and Figure 3 suggests slow decay, but the oscillatory structure makes rigorous conclusions difficult. Determining a global modulus of continuity of $G$ is also open. Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli; Corollary 8 gives $1/c(p)=O(\log\log p)$ pointwise. Closing the gap between the empirical bound $4$ and the asymptotic order $\log\log p$ remains an interesting open problem.
Declaration of competing interests
The author declares that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Research Data Statement
No external datasets were used; all numerical values are reproducible from the explicit formulas and Monte Carlo procedures described in the paper. All empirical values of $c(p)$ in Figures 12 and Table 3 were computed by evaluating the totient ratio $\varphi(p-1)/(p-1)$ from the factorization of p1p-1, followed by the truncated Euler product $P_{J}(p)$ of Proposition 8 with $J=20$; the resulting error is below $10^{-9}$ for every prime $p\geq 3$. The Monte Carlo approximation of $G$ in Figure 1 samples the product distribution of Lemma 10 truncated at the 10,00010{,}000th odd prime $\ell_{N}$. Write $T_{N}:=\sum_{\ell>\ell_{N}}Z_{\ell}$ for the omitted tail and $S_{N}:=\sum_{3\leq\ell\leq\ell_{N}}Z_{\ell}$ for the retained partial sum, so that $X_{N}:=\tfrac{1}{2}e^{-S_{N}}$ approximates $X=\tfrac{1}{2}e^{-S}$ with $S=S_{N}+T_{N}$. The tail mean satisfies $\mathbb{E}[T_{N}]=\sum_{\ell>\ell_{N}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}<10^{-5}$. Since $T_{N}\geq 0$ and $X_{N}\leq\tfrac{1}{2}$, the pointwise coupling bound $0\leq X_{N}-X=X_{N}(1-e^{-T_{N}})\leq T_{N}/2$ holds almost surely, giving $W_{1}(G_{N},G)\leq\tfrac{1}{2}\,\mathbb{E}[T_{N}]<5\times 10^{-6}$ for the Wasserstein-$1$ distance between the truncated and full distribution functions. Markov’s inequality further gives $\mathbb{P}(T_{N}>\varepsilon)<10^{-5}/\varepsilon$ for every $\varepsilon>0$. Figure 3 was produced by evaluating the closed-form Euler product of Proposition 5 over all primes $\ell\leq 10^{7}$. Tables 1 and 2 were computed from the explicit formula $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ of Theorem 9 and the exact factorizations of p1p-1 for the listed moduli, respectively, using only rational arithmetic.